a photogapher wishes to mix 16liters of a 7% acetic solution with a 12% solution to get a 9% solution. how much of 12% solution must be added
Answer by mananth(16946)

a photogapher wishes to mix 16liters of a 7% acetic solution with a 12% solution to get a 9% solution. how much of 12% solution must be added
..
16 liters --------- 7%
x liters----------- 12%
..
mixture
16+x liters -------9%
Equate the concentration
16*0.07+0.12x= (16+x)*0.09
1.12+0.12x=1.44+0.09x
multiply by 100
112+12x=144+9x
3x=32
x=10.66 liters 12% solution